Print Calculator Periodic lable Question 21 of 24
Answer #1

With 37.5% CaCl2 and total mass of mixture 992.7 g

Mass of CaCl2 = 37.5 x 992.7/100 = 372.2625 g

Mass of water = 992.7 - 372.2625 = 620.4375 g